Calculate the correlation coefficient of the following data: x 45 46 46 47 48 49 50 y 44 48 45 48 52 51 49

1
Expert's answer
2021-11-01T16:19:00-0400

"x\u0304 =\t\\frac{45+46+46+47+48+49+50}{7}\t= 47.2857\n\\\\\u0233 =\t\\frac{44+48+45+48+52+51+49}{7}\t= 48.1429"

Now, creating a table



Coefficient of correlation

"(r) =\frac{Σ(x_i - x̄)(y_i - ȳ)}{\sqrt{(Σ( x_i - x̄)^2Σ(y_i - ȳ)^2 )
}}
\\=\frac{23.7143}{\sqrt{(19.4286*50.8571)}}
\\=0.7544"